### v2

# Revisiting Reweighted Risk for Calibration: AURC, Focal Loss, and Inverse Focal Loss

Abstract
Several variants of reweighted risk functionals, such as focal losss, inverse focal loss, and the Area Under the Risk–Coverage Curve (AURC), have been proposed in the literature and claims have been made in relation to their calibration properties. However, focal loss and inverse focal loss propose vastly different weighting schemes. In this paper, we revisit a broad class of weighted risk functions commonly used in deep learning and establish a principled connection between these reweighting schemes and calibration errors. We show that minimizing calibration error is closely linked to the selective classification paradigm and demonstrate that optimizing a regularized variant of the AURC naturally leads to improved calibration. This regularized AURC shares a similar reweighting strategy with inverse focal loss, lending support to the idea that focal loss is less principled when calibration is a desired outcome. Direct AURC optimization offers greater flexibility through the choice of confidence score functions (CSFs). To enable gradient-based optimization, we introduce a differentiable formulation of the regularized AURC using the SoftRank technique. Empirical evaluations demonstrate that our AURC-based loss achieves competitive class-wise calibration performance across a range of datasets and model architectures.
1 Introduction
Over the past decade, advances in deep learning have enabled neural networks to achieve, and in some cases surpass, human-level performance on tasks ranging from image recognition [19] to natural language understanding [1, 33]. However, superior accuracy alone does not guarantee that the predicted confidence scores faithfully reflect true outcome probabilities – a phenomenon known as miscalibration [17]. Miscalibration can lead to critical errors in high-stakes applications where trustworthy uncertainty estimates are essential. To address and quantify this gap, the research community has introduced a range of calibration error metrics [5, 17, 30, 41] to evaluate the discrepancy between predicted confidences and observed outcome frequencies. Recent work has explored both trainable [24, 31, 26, 4, 41] and post-hoc [17, 10, 29, 44, 36] approaches for neural network calibration, achieving impressive reductions in terms of those calibration errors. Among trainable methods, focal loss [32]—which emphasizes hard examples during training—has gained popularity for its calibration benefits. Conversely, inverse focal loss [43], which focuses on the easy examples more, has also been shown to improve calibration in certain settings. This raises a key question: Which weighting scheme is most effective for improving model calibration?
We aim to answer this question by examining the relationship between selective classification [13] and calibration error. We provide a theoretical foundation that links calibration error minimization to the principles of selective classification. Building on this connection, we propose a loss function derived from the Area Under the Risk-Coverage Curve (AURC) [14], originally introduced in the context of selective classification, and demonstrate that it adopts a reweighting scheme similar to that of the inverse focal loss. To enable gradient-based optimization, we leverage the SoftRank [3] method to approximate AURC in a differentiable manner. This adaptation ensures that our AURC-based loss remains differentiable with respect to any choice of confidence score function (CSF).
Our main contributions are summarized as follows:
1 We establish a theoretical connection between calibration error minimization and selective classification by analyzing the structure of calibration error and its relation to the “abstain” strategy at the core of selective classification.
2 Motivated by this connection, we propose a differentiable loss function based on the AURC metric. Although it adopts a reweighting strategy similar to that of inverse focal loss, it offers greater flexibility in the choice of CSFs.
3 We demonstrate through extensive experiments that our proposed AURC-based loss achieves competitive class-wise calibration performance across multiple datasets and network architectures, outperforming existing calibration-aware loss functions.
2 Related work
Post-hoc calibration methods address miscalibration by learning a mapping function that transforms model outputs into calibrated posterior probabilities. Approaches like Platt scaling [40], temperature scaling (TS) [17], and local temperature scaling [10] typically preserve classification accuracy by introducing a scalar temperature parameter. In contrast, other probabilistic methods—including Beta calibration [28], Dirichlet calibration [29], Gaussian processes [44], and the Concrete distribution [11]—may achieve better calibration but often at the cost of reduced classification accuracy. Notably, Dirichlet calibration generalizes Beta calibration to multi-class settings. Ma and Blaschko [36] and Liu et al. [35] proposed post-hoc calibration methods that adjust predicted probabilities based on correctness, using binary classifiers to either distinguish between correct and incorrect predictions or to directly modulate confidence accordingly. However, these approaches often involve a trade-off between calibration and predictive accuracy. We aim to develop a trainable calibration method that improves calibration performance without significantly compromising predictive accuracy.
Trainable calibration methods seek to enhance the calibration of deep neural networks by explicitly modifying the training objective. One class of approaches introduces differentiable surrogate losses that directly approximate the expected calibration error, as demonstrated in [24, 31, 26, 4, 41]. Popordanoska et al. [41] proposed $\mathrm{ECE}^{\mathrm{KDE}}$, a differentiable estimator of calibration error, and introduced the KDE-XE objective, which augments cross-entropy (XE) with an ECE-KDE regularizer to balance accuracy and calibration. Alternatively, several studies replace the standard cross-entropy loss with calibration-aware objectives, including mean squared error loss [23], inverse focal loss [43], and focal loss [32]. Focal loss, in particular, has been empirically shown to improve calibration by mitigating overconfident predictions [37, 6, 25]. Nevertheless, Wang et al. [43] argue that the regularization effect of focal loss may suppress important information about sample difficulty, thereby limiting the effectiveness of subsequent post-hoc calibration techniques. To address this trade-off, several variants of focal loss have been proposed, such as Adaptive Focal Loss [15] and Dual Focal Loss [42]. Distinct from these prior approaches, we propose a loss function based on the selective classification framework. While selective classification is not originally intended for calibration, we show that its underlying principle is fundamentally aligned with minimizing calibration error.
Selective classification augments a standard classifier with a reject option, enabling an explicit trade-off between selective risk and coverage. This technique mitigates selective risk by withholding low-confidence predictions [13, 14, 9, 12]. The Area Under the Risk-Coverage Curve (AURC) and its normalized variant, Excess-AURC (E-AURC) [14], are among the most widely used evaluation metrics for selective classification systems. These metrics compute the risk associated with accepted predictions across different confidence thresholds. Zhou et al. [47] showed that AURC can be interpreted as a weighted risk function, where the weights are determined by the ranking induced by a CSF $g$. This work is motivated by the goal of establishing a theoretical and empirical connection between selective classification and model calibration. We empirically show that minimizing a regularized AURC objective yields strong calibration performance across a range of architectures and datasets. While the underlying weighting scheme is similar to that of the inverse focal loss, our approach offers substantially greater flexibility in the choice of CSFs, making it more adaptable to different modeling scenarios.
3 Preliminaries
A common approach to training a classifier with a softmax output is to learn a function $f:\mathcal{X}\to\Delta^{k}$, which maps an input $\boldsymbol{x}\in\mathbb{R}^{d}$ to a probability distribution over $k$ classes. The output $\boldsymbol{p}=f(\boldsymbol{x})=(p_{1},\dots,p_{k})^{\top}\in\Delta^{k}$ is a vector in the probability simplex, representing the predicted class probabilities. Given an input $\boldsymbol{x}$, let $y^{\prime}\in\{1,\dots,k\}$ denote the true class label, and define its one-hot encoded label $\boldsymbol{y}=(y_{1},\dots,y_{k})^{\top}\in\{0,1\}^{k}$, where $y_{y^{\prime}}=1$ and $y_{j}=0$ for all $j\neq y^{\prime}$.
3.1 Weighted Risk Functions
To train the classifier $f$, one typically minimizes the expected loss over the joint probability distribution $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$:
$\mathcal{R}(f)=\mathbb{E}[\ell(f\left(\boldsymbol{x}\right),\boldsymbol{y})],$ (1)
where $\ell$ is a loss function, such as cross-entropy or squared loss. These conventional loss functions treat all samples equally, which may not be optimal in scenarios with imbalanced data or varying prediction confidence. A widely adopted alternative is the focal loss, a weighted loss function that emphasizes uncertain or misclassified examples.
Definition 3.1 (Focal Loss).
Let $\boldsymbol{p}$ be a predicted probability vector, and $\gamma\geq 0$ a focusing parameter. The focal loss is defined as:
$\ell_{\text{FL}}(\boldsymbol{p},\boldsymbol{y})$ $=-\left(1-p_{y^{\prime}}\right)^{\gamma}\ln(p_{y^{\prime}}),$ (2)
where $p_{y^{\prime}}$ denotes the predicted probability corresponding to the true class label $y^{\prime}$.
Focal loss reduces to cross-entropy when $\gamma=0$. More generally, it can be interpreted as a reweighted version of cross-entropy, where each instance is assigned an importance weight $\boldsymbol{w}(\boldsymbol{p})=(1-\boldsymbol{p})^{\gamma}\in\mathbb{R}^{k}$. The weight depends on both the predictive probability $p_{y^{\prime}}$ and the focusing parameter $\gamma$: samples predicted with high confidence are downweighted, while harder or less confident samples are emphasized when $\gamma\geq 1$.
Definition 3.2 (Inverse Focal Loss).
Let $\boldsymbol{p}$ be a predicted probability vector, $\gamma\geq 0$, and $f$ a model mapping into the probability simplex. The inverse focal loss is defined as:
$\ell_{\text{IFL}}(\boldsymbol{p},\boldsymbol{y})$ $=-\left(1+p_{y^{\prime}}\right)^{\gamma}\ln(p_{y^{\prime}}),$ (3)
where $p_{y^{\prime}}$ denotes the predicted probability for the true class label $y^{\prime}$.
(a) Normalized Weights
(b) Loss with Normalized Weights
Figure 1: Comparison of weight and loss behavior with respect to prediction probability $p$. We normalize the weights in both focal loss and inverse focal loss for fair comparison.
In contrast to focal loss, which downweights confident predictions, inverse focal loss assigns greater weight to high-confidence samples. We analytically derive the gradient of inverse focal loss (App. 7.3) and show empirically that certain values of $\gamma$ may lead to undesirable behavior during training (see Fig. 5 in appendix). Surprisingly, despite their fundamentally opposite reweighting schemes, both strategies have been reported to improve classifier calibration.
3.2 Selective classification
Selective classification pairs a classifier $f$ with a confidence scoring function $g:\mathcal{X}\rightarrow[0,1]$, which estimates the reliability of its predictions. A selective classifier, defined as the pair $(f,g)$, makes a prediction only when the confidence score exceeds a threshold $\tau\in\mathbb{R}$, and abstains otherwise:
$(f,g)(\boldsymbol{x}):=\begin{cases}f(\boldsymbol{x})&\text{if }g(\boldsymbol{x})\geq\tau,\\ \text{"abstain"}&\text{otherwise}.\end{cases}$ (4)
The selective risk w.r.t. the data distribution $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$ is defined as:
$R(f,g):=\frac{\mathbb{E}\left[\ell(f(x),y)\cdot\mathbb{I}\left(g(x)\geq\tau\right)\right]}{\mathbb{E}\left[\mathbb{I}\left(g(x)\geq\tau\right)\right]}.$ (5)
where $\mathbb{I}$ is the indicator function and the denominator represents the Coverage. The threshold $\tau$ determines the trade-off between selective risk and coverage. Rather than evaluating performance at a single threshold, the AURC [14] assesses classifier robustness across all possible thresholds, as induced by the distribution of confidence scores. This metric can equivalently be expressed as a weighted risk function [47], as described below.
Definition 3.3 (AURC).
Let $\boldsymbol{p}=f(\boldsymbol{x})$ be a predicted probability vector produced by the model $f$ with fixed input $\boldsymbol{x}$. Denote with $\boldsymbol{p}^{\prime}=f\left(\boldsymbol{x}^{\prime}\right)$ the output random variable with input random variable $\boldsymbol{x}^{\prime}$. The cumulative distribution function (CDF) of the CSF $g(\boldsymbol{p})$ is defined as:
[…]
Then, the AURC metric is given by:
$\text{AURC}_{g}(f)=-\mathbb{E}\left[\ln\left(1-G(\boldsymbol{p})\right)\cdot\ell(\boldsymbol{p},\boldsymbol{y})\right].$ (7)
This formulation reveals that AURC can be interpreted as a reweighted risk, where each sample is weighted by $w(\boldsymbol{p})=-\ln(1-G(\boldsymbol{p}))$, a function that depends on the CSF. Common choices for the CSFs include the Maximum Softmax Probability (MSP), i.e., $g(\boldsymbol{p})=\max_{i=1}^{k}p_{i}$, or the negative entropy of the logits prior to the softmax layer as described in Table 3.
3.3 Calibration Errors
Below, we provide the definitions of commonly used calibration errors and their empirical estimators.
Definition 3.4 (Calibrated).
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ almost surely based on the input random variable $\mathbf{x}$.
Definition 3.5 ($L_{\rho}$ Calibration Error ($\text{CE}_{\rho}$)).
The $L_{\rho}$ calibration error of a model $f:\mathcal{X}\rightarrow\Delta^{k}$, with $\rho\geq 1$, is defined as:
$\text{CE}_{\rho}(f)=\left(\mathbb{E}\left[\|f(\boldsymbol{x})-\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]\|_{\rho}^{\rho}\right]\right)^{1/\rho},$ (8)
where the expectation is taken with respect to $\boldsymbol{x}$.
When $\rho=1$ and for binarized outcomes, this quantity is referred to as the Expected Calibration Error (ECE) [38, 17], i.e.,
$\operatorname{ECE}\left(f\right)=\mathbb{E}\left[\lvert\max_{c}f_{c}(\boldsymbol{x})-\mathbb{P}(\boldsymbol{y}=\arg\max_{c}f_{c}(\boldsymbol{x})\mid\max_{c}f_{c}(\boldsymbol{x}))\rvert\right].$ (9)
In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown conditional distribution $\mathbb{P}(\boldsymbol{y}\mid f\left(\boldsymbol{x}\right))$. To address this, Naeini et al. [38] proposed an empirical estimator based on a finite dataset. This estimator employs a binning scheme $B$ to approximate the true difference between the predicted confidence and the conditional probability $\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ by computing empirical averages within each bin. We denote this estimator as $\widehat{\text{ECE}}_{B}$.
Definition 3.6 (Binned estimator $\widehat{\text{ECE}}_{B}$).
Let the binning scheme be $B=\{I_{b}\big{|}b=1,\dots,m\}$, the binned estimator is denoted as:
$\widehat{\text{ECE}}_{B}(f)=\sum_{b=1}^{m}\frac{n_{b}}{n}\left|\text{acc}_{b}(f)-\text{conf}_{b}(f)\right|$ (10)
where $n_{b}$ is the number of samples falling into bin $I_{b}$, $n$ is the total number of samples, $\text{conf}_{b}(f)$ is the average predicted confidence in bin $I_{b}$, and $\text{acc}_{b}(f)$ is the accuracy in that bin.
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) [30, 29] measures the expected calibration errors across all classes.
Definition 3.7 (Class-Wise Calibration Error ($\text{cwECE}_{\rho}$)).
The class-wise calibration error is:
$\text{cwECE}_{\rho}(f)=\Big{(}\frac{1}{k}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big{)}^{1/\rho},$ (11)
for $1\leq\rho\in\mathbb{R}$, where $f_{c}(\boldsymbol{x})$ is the predicted probability for class $c$, and $y^{\prime}$ is the true class label.
We follow related works [39, 18] in defining this metric as the average calibration error across classes. Similar to ECE, it can be estimated empirically using a binning scheme $B$, resulting in a bin-based estimator $\widehat{\text{cwECE}}_{B}(f)$.
Definition 3.8 (Binned estimator $\widehat{\text{cwECE}}_{B}$).
Let $m_{c}$ be the number of bins for class $c$ and the binning scheme be $B=\{I_{c,b}\big{|}\ c=1,\dots,k;b=1,\dots,m_{c}\}$, The binned estimator for $\rho=1$ is given by:
$\widehat{\mathrm{cwECE}}_{B}(f)$ $=\frac{1}{k}\sum_{c=1}^{k}\frac{1}{n}\sum_{b=1}^{m_{c}}n_{c,b}\Bigl{|}\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!\!\mathbb{I}(y_{i}^{\prime}=c)}_{\text{acc}_{c,b}}-\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!f_{c}(\boldsymbol{x}_{i})}_{\text{conf}_{c,b}}\Bigr{|}$ (12)
where $n_{c,b}$ is the number of samples in bin $I_{c,b}$, and $n$ is the total number of samples.
4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and class-wise ECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how the calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from the AURC metric in selective classification. The formulation is fully differentiable—enabled by the SoftRank method—and supports arbitrary CSFs.
4.1 Linking Calibration Error Minimization to Selective Classification
Lemma 4.1 (Lower bound of ECE [36, Prop. 2]).
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl{(}\hat{y}_{i}\neq y_{i}^{\prime}\bigr{)}$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top-label ECE computed over all binning schemes $B\in\mathcal{B}$ holds that
$\operatorname{ECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f)\;\geq\;\frac{\widehat{\operatorname{err}}(f)}{k}.$ (13)
In the proof of Prop. 4.1, Ma and Blaschko [36] demonstrate that a calibrated model can be constructed using a binary classifier that separates correct from incorrect predictions. We reinterpret this insight through the lens of selective classification, yielding the following result.
Proposition 4.2 (Optimal Calibration Map for ECE [36, Prop. 3]).
Suppose we are given a CSF $g$ that perfectly distinguishes correct from incorrect predictions via a threshold $\tau$, and a model $f:\mathcal{X}\rightarrow\Delta^{k}$. Then, we can construct a modified model $f^{\prime}$ such that:
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f),$ (14)
with the following adjustment:
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \frac{1}{k}&\text{if }g(\boldsymbol{x})<\tau.\end{cases}$ (15)
Apart from this modification, $f^{\prime}$ is identical to $f$.
This proposition suggests that ECE can be improved by boosting the predicted confidence for correctly classified samples while reducing it for incorrect ones. Notably, assigning a uniform prediction $\frac{1}{k}$ when $g(\boldsymbol{x})<\tau$ mimics the abstention mechanism used in selective classification. In this context, the CSF with a threshold $\tau$ effectively acts as a binary classifier. To further support the connection between calibration errors and selective classification, we next extend Prop. 4.1 to the cwECE and derive a constructive rule for minimizing it.
Proposition 4.3 (Lower Bound of cwECE).
For any class-wise binning scheme $B\in\mathcal{B}$, the class-wise ECE satisfies the lower bound:
$\operatorname{cwECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)>\frac{1}{k}\widehat{\operatorname{err}}(f).$ (16)
The proof is provided in App. 7.2, and we sketch it here. First, we apply Minkowski’s inequality to derive a lower bound on the supremum over all class-wise binning schemes. Next, we show that this supremum is bounded below by $\widehat{\operatorname{err}}(f)/k$, unless $f$ is perfectly accurate. Finally, by invoking Jensen’s inequality, as in Kumar et al. [30], the lower bound follows. This construction naturally leads to a procedure for improving cwECE, analogous to Prop. 4.2.
Proposition 4.4 (Optimal Calibration Map for cwECE).
Let $g$ be a CSF that perfectly separates correct and incorrect predictions via the threshold $\tau$, and let $f:\mathcal{X}\rightarrow\Delta^{k}$. Then, a modified model $f^{\prime}$ can be constructed such that
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)$ (17)
with
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \min\left(\frac{1}{2},\max_{c}f_{c}(\boldsymbol{x})\right)&\text{otherwise}.\end{cases}$ (18)
All other components of $f^{\prime}$ are the same as in $f$.
This construction rule follows directly from the proof of Prop. 4.3. While slightly different from Prop. 4.2, this result similarly suggests that reducing the maximum predicted probability for uncertain samples helps improve cwECE. Together, these propositions build a theoretical foundation connecting selective classification with calibration error minimization: while Prop. 4.2 addresses ECE, Prop. 4.3 and its constructive variant extend this idea to cwECE. We conclude by presenting a result from Gruber and Buettner [16], which synthesizes calibration metrics under a unified theoretical bound.
Proposition 4.5 (Hierarchy Among Calibration Metrics [16, Theorem 3.1]).
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then,
$n^{\frac{1}{\rho}-\frac{1}{2}}\sqrt{\text{BS}(f)}\geq\text{CE}_{\rho}(f)\geq\begin{cases}k\cdot\operatorname{cwECE}_{\rho}(f),\\ \operatorname{ECE}(f),\end{cases}\geq 0,$ (19)
where BS denotes the Brier score.
The proposition formalizes a hierarchy of calibration metrics, showing that $L_{\rho}$ calibration error is upper bounded by a scaled Brier score and lower bounded by both cwECE and ECE. Thus, minimizing ECE or cwECE serves as a practical proxy for improving overall calibration performance.
4.2 Regularized AURC Loss Function
Given the preceding results, reducing calibration error requires lowering the maximum predicted probability for low-confidence samples—a strategy that echoes the "abstain" mechanism central to selective classification. Since the AURC metric originates from the selective classification framework and employs a weighting scheme similar to that of inverse focal loss (see Fig. 1), such reweighting schemes are theoretically well-suited for calibration error minimization. Motivated by this, we propose the regularized AURC (r-AURC) loss to examine whether this reweighting approach can effectively enhance model calibration. An asymptotically unbiased Monte Carlo estimator [47] for the population AURC defined in Eq. (7) is given by
$\widehat{\operatorname{AURC}}_{g}(f)=-\frac{1}{n}\sum_{i=1}^{n}\ln\Big{(}1-\frac{r_{i}}{n+1}\Big{)}\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i}),$ (20)
where $r_{i}:=r(g(\boldsymbol{p}_{i}))$ denotes the rank of the confidence score $g(\boldsymbol{p}_{i})$ among all samples in ascending order. Thus, a larger $r_{i}$ corresponds to a higher confidence. This estimator is consistent and can be directly used as a training loss to minimize the AURC metric. Importantly, since $r_{i}=r(g(\boldsymbol{p}_{i}))$ is a function of the model output $\boldsymbol{p}_{i}$, its gradient $\frac{\partial r_{i}}{\partial\boldsymbol{p}_{i}}$ must be considered in backpropagation. To facilitate end-to-end optimization, we adopt the SoftRank approximation [3], which replaces the hard (non-differentiable) rank with a smooth, differentiable surrogate $\hat{r}_{i}$. This enables standard gradient-based training, as outlined in Algorithm 1.
Algorithm 1 AURC Optimization with $\mathrm{SoftRank}$
1:Training batch $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, model $f_{\theta}$, a CSF $g(\cdot)$, learning rate $\eta$
2:Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}L$
3:Forward Pass:
4:$\boldsymbol{p}_{i}\leftarrow f_{\theta}(\boldsymbol{x}_{i})$triangleright Model prediction
5:$\hat{r}_{i}\leftarrow\mathrm{SoftRank}(g(\boldsymbol{p}_{i}))$
6:$L\leftarrow\frac{1}{n}\sum_{i=1}^{n}\left[-\ln(1-\hat{r}_{i})\cdot\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})\right]$triangleright AURC loss
7:Backward Pass:
8:for each $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})$ in batch do
9:     Compute $\nabla_{\boldsymbol{p}_{i}}L$ via Eq. (41)
10:     $\nabla_{\theta}L\mathrel{+}=(\nabla_{\boldsymbol{p}_{i}}L)^{\top}\nabla_{\theta}\boldsymbol{p}_{i}$
11:end for
This loss function typically relies on sorting based on the confidence score function (CSF), incurring a computational cost of $\mathcal{O}(n\log(n))$, which is generally more efficient than kernel-based calibration error losses that require quadratic time. To better understand the effect of CSF on AURC metric, we define the perfectly aligned CSF.
Definition 4.6 (Perfectly Aligned CSF).
A CSF $g^{\star}$ is said to be perfectly aligned with the loss function $\ell$ if, for any two predictions $\boldsymbol{p}_{i}$, $\boldsymbol{p}_{j}$ with corresponding labels $\boldsymbol{y}_{i}$, $\boldsymbol{y}_{j}$, the following holds:
$g^{\star}(\boldsymbol{p}_{i})\leq g^{\star}(\boldsymbol{p}_{j})\quad\Rightarrow\quad\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})\geq\ell(\boldsymbol{p}_{j},\boldsymbol{y}_{j}).$ (21)
This condition ensures that higher confidence scores correspond to lower losses, meaning that the confidence ranking is consistent with the loss. In practice, this alignment significantly affects the effectiveness of the r-AURC loss, as it governs how the sample weights (via soft rank) interact with the loss values.
Proposition 4.7 (Lower Bound for Empirical AURC).
Let $g^{\star}$ be a CSF that is perfectly aligned with the loss function $\ell$. Let $\mathcal{G}$ denote the set of all valid confidence mappings. For a given model $f$ and the cross-entropy loss, the empirical AURC is minimized at $g^{\star}$, and has the following lower bound:
$\inf_{g\in\mathcal{G}}\widehat{\text{AURC}}_{g}(f)=\widehat{\text{AURC}}_{g^{\star}}(f)\geq\frac{n\log 2}{2(n+1)}\widehat{\operatorname{err}}^{2}(f)+\frac{\log 2}{2(n+1)}\widehat{\operatorname{err}}(f),$ (22)
where $\widehat{\operatorname{err}}(f)$ is the empirical error of $f$ on a finite dataset.
See App. 7.2 for the proof. This proposition indicates that AURC is minimized when the CSFs are perfectly aligned with the loss and when the model achieves high predictive accuracy. In addition, Prop. 4.1 and 4.3 show that lower calibration error also relies on the model being accurate. These insights collectively suggest that both calibration and accuracy are mutually reinforcing goals.
To this end, we consider the following constrained optimization framework to learn a model $f$ that is both accurate and well-calibrated: $f^{*}=\arg\min_{f\in\mathcal{F}}\mathcal{R}(f),\quad\text{subject to}\quad\text{AURC}_{g}(f)\leq B,$ where $\mathcal{R}(f)$ denotes the expected predictive risk, and $B>0$ is a user-defined budget. The corresponding Lagrangian relaxation of this constrained objective is given by:
$f^{*}=\arg\min_{f\in\mathcal{F}}\left((1-\lambda)\,\mathcal{R}(f)+\lambda\,\text{AURC}_{g}(f)\right),$ (23)
where $\lambda\in[0,1]$ controls the trade-off between accuracy and AURC minimization. Note that the behavior of this regularized loss depends on the choice of the CSF $g$ used in the AURC. As we show in later sections, different CSFs may lead to different calibration-performance.
5 Experiments
Datasets and Models. We conduct experiments on image classification datasets including CIFAR-10, CIFAR-100 [27], and Tiny-ImageNet [8]. For Tiny-ImageNet, all images are resized to 224×224 pixels. On CIFAR-10/100, we evaluate models including WideResNet-28x10 [46], PreResNet-56 [21], and ResNet architectures with depths 50 and 110 [20]. For Tiny-ImageNet, we use pre-trained models from the timm library [45]. Each experiment is repeated with five different random seeds, and we report the mean of the resulting performance metrics.
[…]
Evaluation Metrics. We evaluate calibration performance using the binned estimators of ECE and class-wise ECE (cwECE), computed under two binning strategies: equal-width binning (denoted as ECE and cwECE) and adaptive binning [9] (denoted as ECEa and cwECEa). In addition, we consider a kernel density estimation-based calibration error, denoted as $\mathrm{ECE}^{\mathrm{KDE}}$ [41], which provides a continuous, non-binned alternative. To assess whether our proposed r-AURC loss directly improves its intended objective, we also report the AURC metric in all evaluations.
Baselines. We explore the trainable calibration methods including cross-entropy loss (XE), KDE-XE [41], focal loss [37], inverse focal loss [43], and dual focal loss [42]. For focal loss, we consider FL-53 where $\gamma=5$ for $f_{y^{\prime}}(\boldsymbol{x})\in[0,0.2)$ and $\gamma=3$ otherwise. For the latter two loss functions, the $\gamma$ parameter is fine-tuned on the split validation set following the settings in Tao et al. [42]. For inverse focal loss, we note that when $\gamma>3.5$, it may not able to update the model properly due to the positive gradients, as detailed discussed in App. 7.3. For KDE-XE, we set its regularization parameter $\lambda=0.3$ on CIFAR-10, and $\lambda=0.5$ on CIFAR-100, to achieve the best canonical calibration error as suggested in the original paper. We also evaluate post-hoc calibration methods, such as MetaAcc [36] and Temperature Scaling (TS) [17], on models trained with different loss functions. (See App. 7.5 for additional results on those baselines.)
Table 1: Test accuracy (%) and classwise ECEa ($\downarrow$), reported as mean ×$\times 10^{2}$ over 5 seeds.
Accuracy (%) $\uparrow$ cwECEa ($\downarrow$)
Model XE FL-53 Inv.FL r-Ap r-Ae r-Am XE FL-53 Inv.FL r-Ap r-Ae r-Am
CIFAR-10
ResNet50 94.75 94.43 95.00 94.61 94.30 94.63 0.266 0.591 0.287 0.260 0.249 0.259
ResNet110 94.56 94.32 95.05 94.44 94.25 94.53 0.275 0.544 0.294 0.264 0.271 0.269
WRN 95.30 94.03 95.42 95.42 95.17 95.56 0.236 1.059 0.256 0.251 0.244 0.220
PreResNet56 93.39 93.05 93.78 93.36 93.22 93.38 0.308 0.592 0.383 0.324 0.323 0.313
CIFAR-100
ResNet50 77.54 68.08 77.92 77.68 76.30 77.62 0.092 0.239 0.100 0.091 0.091 0.091
ResNet110 77.21 71.13 77.32 77.43 71.87 77.49 0.098 0.178 0.111 0.097 0.131 0.095
WRN 78.68 72.68 78.73 78.73 78.66 78.86 0.094 0.190 0.106 0.092 0.094 0.098
PreResNet56 72.21 66.06 72.63 72.43 72.53 72.34 0.107 0.192 0.113 0.105 0.090 0.106
Tiny-ImageNet
ResNet50 81.84 78.24 79.56 82.11 82.47 82.31 0.126 0.177 0.136 0.127 0.122 0.123
Results. Table 1 presents the test performance in terms of cwECE across different loss functions. Overall, the proposed r-AURC loss consistently outperforms other baselines, while focal loss tends to yield the worst cwECE in most cases. Notably, the performance of r-AURC is sensitive to the choice of CSF; for instance, MSP proves suboptimal on CIFAR-10, underscoring the importance of selecting a suitable CSF to ensure reliable calibration. In Fig. 2, focal loss tends to be overconfident on easy samples, whereas AURC-based losses exhibit better-calibrated behavior. As shown in Table 2, the cwECEa of r-AURC can be further reduced via TS, highlighting the benefit of post-hoc calibration. Additional results in Appendix 7.5 show that MetaAcc can further enhance cwECE—particularly for losses other than focal loss—though this often comes at the cost of reduced predictive accuracy. Regarding ECE, r-AURC achieves the best overall performance on CIFAR-10 across all architectures, and its ECE can likewise be improved through TS. Notably, focal loss performs relatively well in terms of ECE on CIFAR-100, suggesting its utility in top-label settings. Given that focal loss lacks a direct theoretical connection to calibration error minimization, we advocate for the use of inverse focal loss or r-AURC, both of which are explicitly designed to optimize calibration. Our r-AURC loss does not consistently achieve the lowest ECE across all datasets, suggesting a degree of dataset dependency or the need for incorporating a more reliable CSF to fully realize its calibration potential.
Table 2: Mean cwECE ×$\times 10^{2}$ ($\downarrow$) before and after TS. The temperature $T$ selected by minimizing the ECE on the training set is shown in brackets next to each TS result.
Model XE FL-53 Inv. FL r-Ap r-Ae r-Am
CIFAR-10
ResNet50 0.641 0.838 0.805 0.629 0.621 0.644
     +TS 0.456 [1.30] 0.911 [1.02] 0.692 [1.36] 0.438 [1.29] 0.407 [1.28] 0.457 [1.30]
ResNet110 0.670 0.767 0.884 0.691 0.656 0.671
     +TS 0.478 [1.30] 0.870 [1.03] 0.803 [1.38] 0.478 [1.31] 0.424 [1.29] 0.463 [1.30]
WRN 0.623 1.215 0.701 0.605 0.608 0.588
     +TS 0.417 [1.33] 1.093 [0.97] 0.554 [1.37] 0.407 [1.33] 0.393 [1.33] 0.381 [1.34]
PreResNet56 0.639 0.799 0.881 0.606 0.586 0.628
     +TS 0.462 [1.20] 0.634 [0.95] 0.729 [1.30] 0.424 [1.20] 0.428 [1.18] 0.444 [1.21]
CIFAR-100
ResNet50 0.248 0.285 0.268 0.245 0.246 0.246
     +TS 0.189 [1.36] 0.284 [1.02] 0.200 [1.37] 0.192 [1.36] 0.191 [1.30] 0.189 [1.36]
ResNet110 0.259 0.244 0.293 0.257 0.276 0.252
     +TS 0.194 [1.36] 0.255 [1.07] 0.209 [1.38] 0.191 [1.36] 0.224 [1.27] 0.188 [1.36]
WRN 0.235 0.254 0.267 0.238 0.218 0.240
     +TS 0.189 [1.37] 0.285 [1.09] 0.193 [1.38] 0.191 [1.36] 0.204 [1.35] 0.190 [1.36]
PreResNet56 0.271 0.271 0.308 0.264 0.223 0.265
     +TS 0.228 [1.13] 0.273 [0.96] 0.255 [1.18] 0.220 [1.14] 0.196 [1.10] 0.221 [1.14]
Tiny-ImageNet
ResNet50 0.127 0.177 0.136 0.127 0.122 0.123
     +TS 0.152 [1.21] 0.195 [1.07] 0.160 [1.22] 0.153 [1.21] 0.139 [1.17] 0.151 [1.21]
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationOutputGapECE=2.39
(a) r-AURCm
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationOutputGapECE=4.02
(b) Inverse Focal
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationOutputGapECE=3.20
(c) Focal
Figure 2: Reliability diagrams of PreResNet56 on CIFAR10. Top-label ECE reported at the bottom.
6 Conclusion and Limitation
In this work, we establish a theoretical connection between selective classification and calibration error minimization, and propose a differentiable loss function derived from the AURC metric. While this loss adopts a weighting scheme similar to that of inverse focal loss, it provides greater flexibility in the choice of CSFs, which contributes to its superior empirical performance. We introduce a general framework that enables the integration of this metric into a differentiable loss function using any CSF. A limitation of the r-AURC loss is that its performance depends on the quality of the chosen CSF. We therefore expect that incorporating more reliable confidence estimation methods will further improve the calibration performance of our approach.
Acknowledgement
This research received funding from the Flemish Government (AI Research Program) and the Research Foundation - Flanders (FWO) through project number G0G2921N. HZ is supported by the China Scholarship Council. We acknowledge EuroHPC JU for awarding the project ID EHPC-BEN-2025B22-061 and EHPC-BEN-2025B22-037 access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium.