### v3

# Revisiting Reweighted Risk for Calibration: AURC, Focal and Inverse Focal Loss

Abstract
Several variants of reweighted risk functionals, such as focal loss, inverse focal loss, and the Area Under the Risk–Coverage Curve (AURC), have been proposed for improving model calibration, yet their theoretical connections to calibration errors remain unclear. In this paper, we revisit a broad class of weighted risk functions commonly used in deep learning and establish a principled connection between calibration error and selective classification. We show that minimizing calibration error is closely linked to the selective classification paradigm and demonstrate that optimizing selective risk in low-confidence region naturally leads to improved calibration. This loss shares a similar reweighting strategy with dual focal loss but offers greater flexibility through the choice of confidence score functions (CSFs). Our approach uses a bin-based cumulative distribution function (CDF) approximation, enabling efficient gradient-based optimization without requiring expensive sorting and achieving $\mathcal{O}(nK)$ complexity. Empirical evaluations demonstrate that our method achieves competitive calibration performance across a range of datasets and model architectures.
1 Introduction
Over the past decade, advances in deep learning have enabled neural networks to achieve, and in some cases surpass, human-level performance on tasks ranging from image recognition (Haggenmüller et al.,, 2021) to natural language understanding (Achiam et al.,, 2023; Liu et al., 2024a, ). However, achieving high accuracy does not necessarily ensure that the predicted confidence scores align with the true probabilities of outcomes, a phenomenon known as miscalibration (Guo et al.,, 2017). Miscalibration can lead to critical errors in high‐stakes applications where trustworthy uncertainty estimates are essential. To address and quantify this gap, the research community has introduced a range of calibration error metrics (Brier,, 1950; Guo et al.,, 2017; Kumar et al.,, 2019; Popordanoska et al.,, 2022) to evaluate the discrepancy between predicted confidences and observed outcome frequencies. Recent work has explored both trainable (Karandikar et al.,, 2021; Kumar et al.,, 2018; Krishnan and Tickoo,, 2020; Bohdal et al.,, 2023; Popordanoska et al.,, 2022) and post-hoc (Guo et al.,, 2017; Ding et al.,, 2021; Kull et al.,, 2019; Wenger et al.,, 2020; Ma and Blaschko,, 2021) approaches for neural network calibration, achieving impressive reductions in terms of those calibration errors. Among trainable methods, focal loss (Lin et al.,, 2017)—which emphasizes hard examples during training—has gained popularity for its calibration benefits. Conversely, inverse focal loss (Wang et al.,, 2021), which focuses on the easy examples more, has also been shown to improve calibration in certain settings. This raises a key question: Which weighting scheme is most effective for improving model calibration?
We address this question by examining the relationship between selective classification (Geifman and El-Yaniv,, 2017) and calibration error. We provide a theoretical foundation that links calibration error minimization to the principles of selective classification, showing that optimal calibration strategies naturally emerge from selective weighting based on confidence rankings. Building on this connection, we propose a loss function similar to the AURC (Geifman et al.,, 2019), but focused specifically on selective risk in low-confidence regions that require careful calibration. Unlike AURC, our method employs a bin-based CDF approximation that enables efficient gradient-based optimization with $\mathcal{O}(nK)$ complexity. This adaptation ensures computational scalability while maintaining differentiability with respect to any choice of CSFs.
Our main contributions are summarized as follows:
  • We establish a theoretical connection between calibration error minimization and selective classification by analyzing the structure of calibration error and its relation to the “abstain” strategy at the core of selective classification.
  • Motivated by this connection, we propose a computationally efficient, differentiable loss function based on selective classification principles, using a bin-based CDF approximation that avoids expensive sorting operations while maintaining gradient flow for any CSF choice.
  • Through extensive experiments, we show that our proposed method delivers competitive calibration performance across various datasets and network architectures, surpassing the majority of existing calibration-aware loss functions.
2 Related work
Post-hoc calibration methods address miscalibration by learning a mapping function that transforms model outputs into calibrated posterior probabilities. Approaches like Platt scaling (Platt et al.,, 1999), temperature scaling (TS) (Guo et al.,, 2017), and local temperature scaling (Ding et al.,, 2021) typically preserve classification accuracy by introducing a scalar temperature parameter. In contrast, other probabilistic methods—including Beta calibration (Kull et al.,, 2017), Dirichlet calibration (Kull et al.,, 2019), Gaussian processes (Wenger et al.,, 2020), and the Concrete distribution (Esaki et al.,, 2024)—may achieve better calibration but often at the cost of reduced classification accuracy. Notably, Dirichlet calibration generalizes Beta calibration to multi-class settings. Ma and Blaschko, (2021) and Liu et al., 2024b proposed post-hoc calibration methods that adjust predicted probabilities based on correctness, using binary classifiers to either distinguish between correct and incorrect predictions or to directly modulate confidence accordingly. However, these approaches often involve a trade-off between calibration and predictive accuracy. We aim to develop a trainable calibration method that improves calibration performance without significantly compromising predictive accuracy.
Trainable calibration methods seek to enhance the calibration of deep neural networks by explicitly modifying the training objective. One class of approaches introduces differentiable surrogate losses that directly approximate the expected calibration error, as demonstrated in (Karandikar et al.,, 2021; Kumar et al.,, 2018; Krishnan and Tickoo,, 2020; Bohdal et al.,, 2023; Popordanoska et al.,, 2022). Popordanoska et al., (2022) proposed $\mathrm{ECE}^{\mathrm{KDE}}$, a differentiable estimator of calibration error, and introduced the KDE-XE objective, which augments cross-entropy (XE) with an ECE-KDE regularizer to balance accuracy and calibration. Alternatively, several studies replace the standard cross-entropy loss with calibration-aware objectives, including mean squared error loss (Hui,, 2020), inverse focal loss (Wang et al.,, 2021), and focal loss (Lin et al.,, 2017). Focal loss, in particular, has been empirically shown to improve calibration by mitigating overconfident predictions (Mukhoti et al.,, 2020; Charoenphakdee et al.,, 2021; Komisarenko and Kull,, 2024). Nevertheless, Wang et al., (2021) argue that the regularization effect of focal loss might suppress crucial information about sample difficulty, which in turn limits the effectiveness of subsequent post-hoc calibration techniques. To address this trade-off, several variants of focal loss have been proposed, such as Adaptive Focal Loss (Ghosh et al.,, 2022) and Dual Focal Loss (Tao et al.,, 2023). Distinct from these prior approaches, we propose a loss function based on the selective classification framework. While selective classification is not originally intended for calibration, we show that its underlying principle is fundamentally aligned with minimizing calibration error.
Selective classification augments a standard classifier with a reject option, enabling an explicit trade-off between selective risk and coverage. This technique mitigates selective risk by withholding low-confidence predictions (Geifman and El-Yaniv,, 2017; Geifman et al.,, 2019; Ding et al.,, 2020; Galil et al.,, 2023). The Area Under the Risk-Coverage Curve (AURC) and its normalized variant, Excess-AURC (E-AURC) (Geifman et al.,, 2019), are among the most widely used evaluation metrics for selective classification systems. These metrics compute the risk associated with accepted predictions across different confidence thresholds. Zhou et al., (2025) showed that AURC can be interpreted as a weighted risk function, where the weights are determined by the ranking induced by a CSF. Our work is motivated by the goal of establishing a theoretical and empirical connection between selective classification and model calibration. We propose a differentiable loss function that emphasize the selective risks in low-confidence region where calibration is most critical. Unlike AURC computation that requires expensive sorting operations, our approach uses a bin-based CDF approximation to achieve computational efficiency while maintaining differentiability, making it practical for large-scale deep learning applications.
3 Preliminaries
A common approach to training a classifier with a softmax output is to learn a function $f:\mathcal{X}\to\Delta^{k}$, which maps an input $\boldsymbol{x}\in\mathbb{R}^{d}$ to a probability distribution over $k$ classes. The output $\boldsymbol{p}=f(\boldsymbol{x})=(p_{1},\dots,p_{k})^{\top}\in\Delta^{k}$ is a vector in the probability simplex, representing the predicted class probabilities. Given an input $\boldsymbol{x}$, let $y^{\prime}\in\{1,\dots,k\}$ denote the true class label, and define its one-hot encoded label $\boldsymbol{y}=(y_{1},\dots,y_{k})^{\top}\in\{0,1\}^{k}$, where $y_{y^{\prime}}=1$ and $y_{j}=0$ for all $j\neq y^{\prime}$.
3.1 Weighted Risk Functions
To train the classifier $f$, one typically minimizes the expected loss over the joint probability distribution $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$:
$\mathcal{R}(f)=\mathbb{E}[\ell(f\left(\boldsymbol{x}\right),\boldsymbol{y})],$ (1)
where $\ell$ is a loss function, such as cross-entropy or squared loss. These conventional loss functions treat all samples equally, which may not be optimal in scenarios with imbalanced data or varying prediction confidence. A widely adopted alternative is the focal loss, a weighted loss function that emphasizes uncertain or misclassified examples.
Definition 3.1 (Focal Loss).
Let $\boldsymbol{p}$ be a predicted probability vector, and $\gamma\geq 0$ a focusing parameter. The focal loss is defined as:
$\ell_{\text{FL}}(\boldsymbol{p},\boldsymbol{y})$ $=-\left(1-p_{y^{\prime}}\right)^{\gamma}\ln(p_{y^{\prime}}),$ (2)
where $p_{y^{\prime}}$ denotes the predicted probability corresponding to the true class label $y^{\prime}$.
Focal loss reduces to cross-entropy when $\gamma=0$. More generally, it can be interpreted as a reweighted version of cross-entropy, where each instance is assigned an importance weight $\boldsymbol{w}(\boldsymbol{p})=(1-\boldsymbol{p})^{\gamma}\in\mathbb{R}^{k}$. The weight depends on both the predictive probability $p_{y^{\prime}}$ and the focusing parameter $\gamma$: samples predicted with high confidence are downweighted, while harder or low confident samples are emphasized when $\gamma\geq 1$.
Definition 3.2 (Inverse Focal Loss).
Let $\boldsymbol{p}$ be a predicted probability vector, $\gamma\geq 0$, and $f$ a model mapping into the probability simplex. The inverse focal loss is defined as:
$\ell_{\text{IFL}}(\boldsymbol{p},\boldsymbol{y})$ $=-\left(1+p_{y^{\prime}}\right)^{\gamma}\ln(p_{y^{\prime}}),$ (3)
where $p_{y^{\prime}}$ is the predicted probability for true class $y^{\prime}$.
(a) Normalized Weights
(b) Loss
Figure 1: Comparison of weight and loss (with normalized weights) behavior with respect to prediction probability pp. We normalize the weights in both focal loss and inverse focal loss for fair comparison.
In contrast to focal loss, which downweights confident predictions, inverse focal loss assigns greater weight to high-confidence samples. We analyze the gradient of inverse focal loss and show empirically that certain values of $\gamma$ may lead to undesirable behavior during training (see Fig. 4 in Appendix C). Surprisingly, despite their fundamentally opposite reweighting schemes, both strategies have been reported to improve classifier calibration.
3.2 Selective classification
Selective classification pairs a classifier $f$ with a confidence scoring function $g:\mathcal{X}\rightarrow\mathbb{R}$, which estimates the reliability of its predictions. A selective classifier, defined as the pair $(f,g)$, makes a prediction only when the confidence score exceeds a threshold $\tau\in\mathbb{R}$, and abstains otherwise:
$(f,g,\tau)(\boldsymbol{x}):=\begin{cases}f(\boldsymbol{x})&\text{if }g(\boldsymbol{x})\geq\tau,\\ \text{"abstain"}&\text{otherwise}.\end{cases}$ (4)
The selective risk w.r.t. the data distribution $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$ is defined as:
$\mathcal{R}(f,g,\tau):=\frac{\mathbb{E}\left[\ell(f(\boldsymbol{x}),\boldsymbol{y})\cdot\mathbb{I}\left(g(\boldsymbol{x})\geq\tau\right)\right]}{\mathbb{E}\left[\mathbb{I}\left(g(\boldsymbol{x})\geq\tau\right)\right]}.$ (5)
where $\mathbb{I}$ is the indicator function and the denominator represents the Coverage. The threshold $\tau$ determines the trade-off between selective risk and coverage.
Rather than evaluating performance at a single threshold, the AURC (Geifman et al.,, 2019) assesses classifier robustness across all possible thresholds, as induced by the distribution of confidence scores. This metric can equivalently be expressed as a weighted risk function (Zhou et al.,, 2025), as described below.
Definition 3.3 (AURC).
Let $f(\boldsymbol{x})$ be a predicted probability vector for fixed input $\boldsymbol{x}$. Denote $\boldsymbol{x}^{\prime}$ as random input variable, the cumulative distribution function (CDF) of the CSF $g(\boldsymbol{x})$ is defined as:
$G(\boldsymbol{x})=\Pr\left(g(\boldsymbol{x}^{\prime})\leq g(\boldsymbol{x})\right)=\int\mathbb{I}\left[g(\boldsymbol{x}^{\prime})\leq g(\boldsymbol{x})\right]\,d\mathbb{P}(\boldsymbol{x}^{\prime}).$ (6)
Then, the AURC metric is given by:
$\text{AURC}(f)=-\mathbb{E}\left[\ln\left(1-G(\boldsymbol{x})\right)\cdot\ell(f(\boldsymbol{x}),\boldsymbol{y})\right].$
This formulation shows that AURC can be interpreted as a reweighted risk, utilizing a weighting scheme similar to that of the inverse focal loss (see Fig. 1). In this scheme, each sample is weighted by a function that depends on the CSF. Common choices for the CSFs include Maximum Softmax Probability (MSP) (Hendrycks and Gimpel,, 2022), Negative Entropy (Liu et al.,, 2020), and Softmax Margin (Belghazi and Lopez-Paz,, 2021) as described in Appendix Table 1.
3.3 Calibration Errors
Below, we provide the definitions of commonly used calibration errors and their empirical estimators.
Definition 3.4 (Calibrated).
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ almost surely based on the input random variable $\boldsymbol{x}$.
Definition 3.5 ($L_{\rho}$ Calibration Error ($\text{CE}_{\rho}$)).
The $L_{\rho}$ calibration error of a model $f:\mathcal{X}\rightarrow\Delta^{k}$, with $\rho\geq 1$, is defined as:
$\text{CE}_{\rho}(f)=\left(\mathbb{E}\left[\|f(\boldsymbol{x})-\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]\|_{\rho}^{\rho}\right]\right)^{1/\rho},$
where the expectation is taken with respect to $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$.
When $\rho=1$ and for binarized outcomes, this quantity is referred to as the Expected Calibration Error (ECE) (Naeini et al.,, 2015; Guo et al.,, 2017). In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown conditional distribution $\mathbb{P}(\boldsymbol{y}\mid f\left(\boldsymbol{x}\right))$. To address this, Naeini et al., (2015) proposed an empirical estimator that uses a binning scheme $B$ to approximate the calibration error, denoted by $\widehat{\text{ECE}}_{B}$.
Definition 3.6 (Binned estimator $\widehat{\text{ECE}}_{B}$).
Let the binning scheme be $B=\{I_{b}\big|b=1,\dots,m\}$, the binned estimator is denoted as:
$\widehat{\text{ECE}}_{B}(f)=\sum_{b=1}^{m}\frac{n_{b}}{n}\left|\text{acc}_{b}(f)-\text{conf}_{b}(f)\right|$
where $n_{b}$ is the number of samples falling into bin $I_{b}$, $n$ is the total number of samples, $\text{conf}_{b}(f)$ is the average predicted confidence in bin $I_{b}$, and $\text{acc}_{b}(f)$ is the accuracy in that bin.
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) (Kumar et al.,, 2019; Kull et al.,, 2019) measures the expected calibration errors across all classes.
Definition 3.7 (Class-Wise Calibration Error ($\text{cwECE}_{\rho}$)).
The class-wise calibration error is:
$\text{cwECE}_{\rho}(f)=\Big(\frac{1}{k}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big)^{1/\rho}$
for $1\leq\rho\in\mathbb{R}$, where $f_{c}(\boldsymbol{x})$ is the predicted probability for class $c$, and $y^{\prime}$ is the true class label.
We follow related works (Panchenko et al.,, 2022; Gupta and Ramdas,, 2022) in defining this metric as the average calibration error across classes. Similar to ECE, it can be estimated empirically using a binning scheme $B$, resulting in a bin-based estimator $\widehat{\text{cwECE}}_{B}(f)$ (see Appendix A.1).
4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and cwECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from selective classification. The formulation is fully differentiable—enabled by a bin-based CDF approximation and supports arbitrary CSFs.
4.1 Linking Calibration Error Minimization to Selective Classification
Lemma 4.1 (Lower bound of ECE (Ma and Blaschko,, 2021, Prop. 2)).
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl(\hat{y}_{i}\neq y_{i}^{\prime}\bigr)$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top‑label ECE computed over all binning schemes $B\in\mathcal{B}$ holds that
$\operatorname{ECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f)\;\geq\;\frac{\widehat{\operatorname{err}}(f)}{k}.$ (7)
In the proof of Prop. 4.1, Ma and Blaschko, (2021) demonstrate that a calibrated model can be constructed using a binary classifier that separates correct from incorrect predictions. We reinterpret this insight through the lens of selective classification, yielding the following result.
Proposition 4.2 (Optimal Calibration Map for ECE (Ma and Blaschko,, 2021, Prop. 3)).
Suppose we are given a CSF $g$ that perfectly distinguishes correct from incorrect predictions via a threshold $\tau$, and a model $f:\mathcal{X}\rightarrow\Delta^{k}$. Then, we can construct a modified model $f^{\prime}$ such that:
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{ECE}}_{B}(f),$
with the following adjustment:
$\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \frac{1}{k}&\text{if }g(\boldsymbol{x})<\tau.\end{cases}$
Apart from this modification, $f^{\prime}$ is identical to $f$.
This proposition suggests that ECE can be improved by boosting the predicted confidence for correctly classified samples while reducing it for incorrect ones. Notably, assigning a uniform prediction $\frac{1}{k}$ when $g(\boldsymbol{x})<\tau$ mimics the abstention mechanism in selective classification. In this context, the CSF with a threshold $\tau$ effectively acts as a binary classifier. To further support the connection between calibration errors and selective classification, we next extend Prop. 4.1 to the cwECE and derive a constructive rule for minimizing it.
Proposition 4.3 (Lower Bound of cwECE).
For any class-wise binning scheme $B\in\mathcal{B}$, the class-wise ECE satisfies the lower bound:
$\operatorname{cwECE}(f)\geq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)>\frac{1}{k}\widehat{\operatorname{err}}(f).$
Proof.
See Appendix B.1. ∎
This construction naturally leads to a procedure for improving cwECE, analogous to Prop. 4.2.
Proposition 4.4 (Optimal Calibration Map for cwECE).
Let $g$ be a CSF that perfectly separates correct and incorrect predictions via the threshold $\tau$, and let $f:\mathcal{X}\rightarrow\Delta^{k}$. Then, a modified model $f^{\prime}$ can be constructed such that
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)$
$\text{with }\max_{c}f^{\prime}_{c}(\boldsymbol{x})=\begin{cases}1&\text{if }g(\boldsymbol{x})\geq\tau,\\ \min\left(\frac{1}{2},\max_{c}f_{c}(\boldsymbol{x})\right)&\text{otherwise}.\end{cases}$
All other components of $f^{\prime}$ are the same as in $f$.
This construction rule follows directly from the proof of Prop. 4.3. While slightly different from Prop. 4.2, this result similarly suggests that reducing the maximum predicted probability for uncertain samples helps improve cwECE. Together, these propositions build a theoretical foundation connecting selective classification with calibration error minimization: while Prop. 4.2 addresses ECE, Prop. 4.3 and its constructive variant extend this idea to cwECE. We conclude by presenting a result from Gruber and Buettner, (2022), which synthesizes calibration metrics under a unified theoretical bound.
Proposition 4.5 (Hierarchy Among Calibration Metrics (Gruber and Buettner,, 2022, Theorem 3.1)).
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then:
$n^{\frac{1}{\rho}-\frac{1}{2}}\sqrt{\text{BS}(f)}\geq\text{CE}_{\rho}(f)\geq\begin{cases}k\cdot\operatorname{cwECE}_{\rho}(f),\\ \operatorname{ECE}(f),\end{cases}\geq 0$
where BS denotes the Brier score.
The proposition formalizes a hierarchy of calibration metrics, showing that $L_{\rho}$ calibration error is upper bounded by a scaled Brier score and lower bounded by both cwECE and ECE. Thus, minimizing ECE or cwECE serves as a practical proxy for improving overall calibration performance.
4.2 Selective Risk Loss
Given the preceding results from Prop. 4.2 and Prop. 4.4, reducing calibration error requires lowering the maximum predicted probability for incorrectly predicted samples and increasing it for correctly predicted ones. The latter is typically handled by the loss function during training. However, the former has received less attention in trainable calibration methods, and this is the issue we aim to address in this paper from the perspective of selective classification. Motivated by this observation, we propose a loss function similar to AURC, but excludes selective risk over the high-confidence region as these samples are well-calibrated by the loss function during training. Let us denote the low confidence region $\mathcal{X}_{\tau}$ as $\mathcal{X}_{\tau}=\{\boldsymbol{x}\mid g(\boldsymbol{x})<\tau\}.$ Then, we focus on the selective risk corresponds to the samples $\tilde{x}$ drawn from the distribution
$\tilde{\boldsymbol{x}}\sim P(\boldsymbol{x}\mid g(\boldsymbol{x})<\tau):=P_{\mathcal{X}_{\tau}},$ (8)
which leads to the selective area under risk converage curve ($\text{AU}_{P_{\mathcal{X}_{\tau}}}$):
$\text{AU}_{P_{\mathcal{X}_{\tau}}}(f)=\mathbb{E}_{\tilde{\boldsymbol{x}}\sim P_{\mathcal{X}_{\tau}}}\left[\frac{\mathbb{E}_{P(x,y)}\left[\ell(f(\boldsymbol{x}),\boldsymbol{y})\mathbb{I}[g(\boldsymbol{x})\geq g(\tilde{\boldsymbol{x}})]\right]}{\mathbb{E}_{\boldsymbol{x}^{\prime}\sim P_{x}}\mathbb{I}[g(\boldsymbol{x}^{\prime})\geq g(\tilde{\boldsymbol{x}})]}\right$]. (9)
Noticing that the expectation in the numerator can be swapped with the expectation outside, the equation above can then be written as:
$\text{AU}_{P_{\mathcal{X}_{\tau}}}(f)=\mathbb{E}_{\text{P}(\boldsymbol{x},\boldsymbol{y})}\left[\beta(\boldsymbol{x})\ell\left(f\left(\boldsymbol{x}\right),\boldsymbol{y}\right)\right]$ (10)
where
[…]
Proposition 4.6 (An Equivalent Expression for $\beta(x)$).
Given the CDF of the CSF $g$ with its inverse denoted as $g^{-1}$, the expression for $\beta(\boldsymbol{x})$ from Eq. (11) is equivalent to
$\beta(\boldsymbol{x})=\begin{cases}-\ln(1-G(g^{-1}(\tau)))&\text{if }\boldsymbol{x}\in\mathcal{X}\setminus\mathcal{X}_{\tau},\\ -\ln(1-G(\boldsymbol{x}))&\text{if }\boldsymbol{x}\in\mathcal{X}_{\tau}.\end{cases}$ (12)
Proof.
See Appendix B.2. ∎
[Bin-based CDF Approximation.] Let $G(\boldsymbol{x})$ denote the CDF of the CSF, as defined in (6), under the data distribution. We approximate $G$ using a binning-based approach. Given bin edges $\mathcal{E}=\{e_{0},e_{1},\ldots,e_{K}\}$ with $e_{0}<\cdots<e_{K}$, the corresponding bin centers are defined as:
$c_{k}=\frac{e_{k}+e_{k+1}}{2},\quad k=0,\dots,K-1.$ (13)
For a set of confidence scores $\{g(\boldsymbol{x}_{i})\}_{i=1}^{n}$, let $h_{k}$ denote the count in each bin:
$h_{k}=\sum_{i=1}^{n}\mathbb{I}[e_{k}\leq g(\boldsymbol{x}_{i})<e_{k+1}].$ (14)
To maintain differentiability for backpropagation, we approximate the CDF using a smoothed step function based on the sigmoid function. The estimated CDF at any point $\boldsymbol{x}$ is:
$\widehat{G}_{\nu}(\boldsymbol{x})=\frac{1}{n}\sum_{k=0}^{K-1}h_{k}\,\sigma\!\left(\frac{\boldsymbol{x}-c_{k}}{\nu}\right),\quad\sigma(t)=\frac{1}{1+e^{-t}},$ (15)
where $\nu>0$ is the smoothing parameter. A detailed discussion of this parameter is provided in Appendix E.2.
[Finite-Sample Estimator.] Given a threshold $\tau$ defining the low-confidence region $\mathcal{X}_{\tau}$, we estimate the sample weights in Eq. (12) as:
$\widehat{\beta}_{i}=\begin{cases}-\ln\!\big(1-\widehat{G}_{\nu}(g^{-1}(\tau))\big),&g(\boldsymbol{x}_{i})\geq\tau,\\[4.30554pt] -\ln\!\big(1-\widehat{G}_{\nu}(\boldsymbol{x}_{i})\big),&g(\boldsymbol{x}_{i})<\tau.\end{cases}$ (16)
This weighting assigns uniform weights to high-confidence samples and rank-based weights to low-confidence samples. Substituting (16) into Eq. (10) yields the differentiable finite-sample estimator:
$\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}(f)=\frac{1}{n}\sum_{i=1}^{n}\widehat{\beta}_{i}\cdot\ell\big(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\big).$ (17)
Algorithm 1 Optimization with $\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}$ loss
1:Input: training data $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$, model $f_{\theta}$, CSF $g(\cdot)$, bin edges $\mathcal{E}$, smoothing parameter $\nu$, low-confidence threshold $\tau$, learning rate $\eta$
2:Forward Pass:
3:  $\boldsymbol{p}_{i}\leftarrow f_{\theta}(\boldsymbol{x}_{i})$
4:  Update $\widehat{G}_{\nu}(\boldsymbol{x})$ with $g(\boldsymbol{p}_{i})$ via Eq. (15)
5:  Compute $\widehat{\beta}_{i}$ via Eq. (16)
6:  $\widehat{\mathrm{AU}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\widehat{\beta}_{i}\cdot\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})$ {Eq. (17)}
7:Backward Pass:
8:for each $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})$ in batch do
9:  Compute $\nabla_{\boldsymbol{p}_{i}}\widehat{\mathrm{AU}}$ via Eq.(27)
10:  $\nabla_{\theta}\widehat{\mathrm{AU}}\mathrel{+}=(\nabla_{\boldsymbol{p}_{i}}\widehat{\mathrm{AU}})^{\top}\nabla_{\theta}\boldsymbol{p}_{i}$
11:end for
12:Parameter Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\widehat{\mathrm{AU}}$
Unlike the original AURC computation, which requires sorting all $n$ samples based on the CSF and thus incurs a complexity of $\mathcal{O}(n\log n)$, the proposed bin-based loss does not require any sorting. Instead, it approximates the CDF of confidence scores using $K$ bins, resulting in a computational cost of $\mathcal{O}(nK)$, which is typically lower than that of AURC and generally more efficient than kernel-based calibration error losses, which require quadratic time complexity. As with AURC, the behavior of this loss depends on the choice of CSF $g$, and, as we show in later sections, different CSFs may lead to different calibration performance.
(a) CIFAR-10
(b) CIFAR-100
(c) Tiny-ImageNet
Figure 2: Test ECE of $\widehat{\mathrm{AU}}_{P_{\mathcal{X}_{\tau}}}$ loss with different CSFs, reported as mean $\times 10^{2}$.
5 Experiments
Table 1: Test accuracy and ECE ($\downarrow$), reported as mean $\times 10^{2}$.
Model XE FL-53 Inv.FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ XE FL-53 Inv.FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
CIFAR-10    Accuracy ($\uparrow$) ECE ($\downarrow$)
[…]
ResNet110 94.47 93.80 93.92 93.72 94.68 93.05 93.39 3.04 4.44 4.80 2.08 2.80 2.44 1.76
[…]
PreResNet56 93.97 92.97 94.14 93.72 94.09 92.63 92.79 2.62 5.50 3.89 2.07 2.57 1.75 0.88
CIFAR-100    Accuracy ($\uparrow$) ECE ($\downarrow$)
[…]
ResNet110 75.27 77.48 77.36 76.30 78.27 74.33 74.11 7.29 4.57 10.90 4.60 9.09 5.23 2.73
[…]
PreResNet56 74.28 72.93 74.18 72.96 74.69 71.91 73.89 7.10 2.44 9.81 0.96 7.42 4.88 1.81
Tiny-ImageNet Accuracy ($\uparrow$) ECE ($\downarrow$)
[…]
Swin-Tiny 85.36 84.31 85.64 85.42 85.36 82.35 84.38 1.62 7.69 3.00 1.60 1.71 2.02 1.45
Dataset Model XE FL-53 Inv. FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
[…]
Table 2: Mean cwECE $\times 10^{2}$ ($\downarrow$) for original models and after applying post-hoc calibration methods including TS. Temperature values are included in parentheses, and the temperature $T$ is selected by minimizing ECE on the training set.
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=5.83
(a) FL-53
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=3.92
(b) Inverse Focal
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=1.92
(c) Dual Focal
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=2.38
(d) KDE-XE
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=2.98
(e) AURC
0.00.20.40.60.81.00.00.20.40.60.81.0ConfidenceAccuracyPerfect CalibrationECE=0.92
(f) $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
Figure 3: Reliability diagrams of PreResNet56 on CIFAR-10. Top-label ECE (%) appears below each panel.
Datasets and Models. We conduct experiments on image classification datasets including CIFAR10/100 (Krizhevsky,, 2009), and Tiny-ImageNet (Deng et al.,, 2009). On CIFAR-10/100, we evaluate models including WideResNet-28x10 (Zagoruyko and Komodakis,, 2016), PreResNet-56 (He et al., 2016b, ), and ResNet architectures with depths 50 and 110 (He et al., 2016a, ). For Tiny-ImageNet, we use two pretrained transformer-based models from the timm library (Wightman,, 2019): ViT-Small (Dosovitskiy et al.,, 2020) and Swin-Tiny (Liu et al.,, 2021). Each experiment is repeated with three different seeds, and we report the mean performance across these runs.
Experimental setup. For experiments on CIFAR-10/100, we follow the training setup described in Xia et al., (2022). For transformer-based models on Tiny-ImageNet, we adopt the training strategy from Yoshioka, (2024). See Appendix D.1 for details.
Evaluation Metrics. We evaluate calibration performance on the test set using binned estimators of ECE and cwECE, computed with two binning strategies: equal-width binning (denoted as ECE and cwECE) and adaptive binning (Ding et al.,, 2020) (denoted as ECEa and cwECEa).
Baselines. We explore the trainable calibration methods including cross-entropy(XE), KDE-XE, focal, inverse focal, AURC, and dual focal loss. For focal loss, we consider FL-53 where $\gamma=5$ for $f_{y^{\prime}}(\boldsymbol{x})\in[0,0.2)$ and $\gamma=3$ otherwise. For inverse focal loss, we note that when $\gamma>3.5$, it may not able to update the model properly due to the positive gradients, as discussed in Appendix C.2. For dual focal loss, the implementation [^1] https://github.com/Linwei94/ICML2023-DualFocalLoss acts in the opposite direction of the intended principle but follows the concept of selective classification, as discussed in Appendix C.3. We additionally evaluate post-hoc calibration on models trained with the XE loss, considering both accuracy-preserving methods i.e. TS and methods that may alter accuracy, including MetaAcc, MetaMis [^2] https://github.com/maxc01/metacal (Ma and Blaschko,, 2021).
Model selection. For inverse focal and dual focal losses, the $\gamma$ parameter is fine-tuned on a split validation set following Tao et al., (2023). For the $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ loss in Eq. (17), the smooth parameter $\mu$ and choice of CSF are selected using the same strategy. See Appendix E for details.
Influence of the CSFs. We evaluate the $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ loss across different CSFs, including MSP, Negative Entropy, and Softmax Margin. As illustrated in Figure 2, the calibration performance of $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ exhibits noticeable sensitivity to the choice of CSF in most scenarios. For instance, Softmax Margin achieves the best ECE on CIFAR-100 except for the PreResNet56 model. On Tiny-ImageNet, Softmax Margin gives the best performance for the ViT-Small, whereas Negative Entropy yields superior results for Swin-Tiny model.
Results for trainable calibration methods. Table 1 presents the test performance in terms of ECE across various loss functions. Overall, the proposed $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ loss demonstrates good calibration performance compared to baseline methods, while focal loss and inverse focal loss are more variable, depending on both the model architecture and dataset. As visualized in Fig. 3, focal loss tends to produce overconfident predictions on samples, whereas inverse focal loss and AURC exhibit underconfidence. In contrast, $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ yields more balanced and better-calibrated confidence scores. We also observe that Dual Focal loss exhibits strong performance across all datasets and models; however, this may be attributed to its implementation, which operates in the opposite direction of the stated objective and resembles the principal of selective classification, as discussed in Appendix C. Moreover, results in Appendix Table 5 shows that the top-label calibration performance can be further improved via TS, highlighting the benefit of post-hoc methods. For non–accuracy-preserving methods i.e. MetaAcc, MetaMis, they yield additional reductions in both cwECE and ECE (see Appendix Table 4) but often at the expense of predictive accuracy.
Table 5 provides an evaluation of cwECE and accuracy across different models and datasets. The $\mathbf{\text{AU}_{P{\mathcal{X}_{\tau}}}}$ loss consistently demonstrates good calibration performance, achieving the best cwECE in the majority of cases. Furthermore, its calibration performance can be further enhanced with TS on the CIFAR-10/100 datasets. In contrast, FL-53 performs poorly relative to other trainable methods on both CIFAR-10 and Tiny-ImageNet, indicating its limited effectiveness in class-wise calibration settings.
6 Conclusion and Limitation
In this work, we establish a theoretical connection between selective classification and calibration error minimization, offering insights into why certain reweighting schemes effectively improve calibration. Building on this foundation, we propose a differentiable loss function derived from selective classification principles, which targets the selective risk in low-confidence regions where calibration improvements are most crucial. Our approach utilizes a bin-based CDF approximation, ensuring computational efficiency while maintaining differentiability for any choice of CSFs.
Interestingly, we find that the dual focal loss implementation implicitly follows similar principles by down-weighting hard examples, although this connection was not explicitly made in their work. In contrast, our method is directly grounded in principled theoretical foundations, contributing to its superior empirical performance across diverse datasets and architectures. A limitation of our approach is its reliance on the quality of the chosen CSF. We therefore anticipate that incorporating more reliable confidence estimation will further enhance calibration performance.
Acknowledgement
This research received funding from the Flemish Government (AI Research Program) and the Research Foundation - Flanders (FWO) through project number G0G2921N. HZ is supported by the China Scholarship Council. We acknowledge EuroHPC JU for awarding the project ID EHPC-BEN-2025B22-061 and EHPC-BEN-2025B22-037 access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium.
Supplementary Materials
Appendix A Definitions
Definition A.1 (Binned estimator $\widehat{\text{cwECE}}_{B}$).
Let $m_{c}$ be the number of bins for class $c$ and the binning scheme be $B=\{I_{c,b}\big|\ c=1,\dots,k;b=1,\dots,m_{c}\}$, The binned estimator for $\rho=1$ is given by:
$\widehat{\mathrm{cwECE}}_{B}(f)$ $=\frac{1}{k}\sum_{c=1}^{k}\frac{1}{n}\sum_{b=1}^{m_{c}}n_{c,b}\Bigl|\text{acc}_{c,b}-\text{conf}_{c,b}\Bigr|$ (18)
where
$\text{acc}_{c,b}=\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!\!\mathbb{I}(y_{i}^{\prime}=c),\quad\text{conf}_{c,b}=\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!f_{c}(\boldsymbol{x}_{i}),$
$n_{c,b}$ is the number of samples in bin $I_{c,b}$, and $n$ is the total number of samples.
Definition A.2 (Brier Score (BS)).
The Brier Score of a model $f:\mathcal{X}\rightarrow\Delta^{k}$ is defined as:
$\text{BS}(f)=\mathbb{E}\left[\|f(\boldsymbol{x})-\boldsymbol{y}\|_{2}^{2}\right]$ (19)
where $f(\boldsymbol{x})$ is the predicted probabilities and $\boldsymbol{y}$ is the one-hot encoded label.
CSFs. The CSFs are generally defined as functions of the predicted probabilities $\boldsymbol{p}$, which are the outputs by passing the logits $\boldsymbol{z}$ produced by the model for the input $\boldsymbol{x}$ through the softmax function $\sigma(\cdot)$, expressed as $\boldsymbol{p}=\sigma(\boldsymbol{z})\in\mathbb{R}^{k}$. The specific forms of these CSFs are outlined as follows:
Method Equation
MSP $g(\boldsymbol{p})=\max_{i=1}^{k}p_{i}$
Softmax Margin
$g(\boldsymbol{p})=p_{i}-\max_{j\neq i}p_{j}$
with $i=\arg\max_{i=1}^{k}p_{i}$
Negative Entropy $g(\boldsymbol{p})=\sum_{i=1}^{k}\boldsymbol{z}_{i}\log\boldsymbol{z}_{i}$
Table 1: Commonly Used CSFs
Appendix B MISSING PROOFS
B.1 Proof of Prop. 4.3
Proof.
For any class‑wise binning scheme $B=\{I_{c,b}\big|\ c=1,\dots,k;b=1,\dots,m_{c}\}$, let $m_{c}$ be the number of bins for class $c$ and $n_{c,b}$ be the number of samples falling into $I_{c,b}$, then by Minkowski inequality, we have
$\widehat{\mathrm{cwECE}}_{B}(f)$ $=\frac{1}{k}\sum_{c=1}^{k}\frac{1}{n}\sum_{b=1}^{m_{c}}n_{c,b}\Bigl|\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!\!\mathbb{I}(y_{i}^{\prime}=c)}_{\text{acc}_{c,b}}-\underbrace{\frac{1}{n_{c,b}}\!\!\sum_{i\in I_{c,b}}\!f_{c}(\boldsymbol{x}_{i})}_{\text{conf}_{c,b}}\Bigr|$
$\leq\frac{1}{nk}\sum_{c=1}^{k}\sum_{b=1}^{m_{c}}\sum_{i\in I_{c,b}}\Bigl|\mathbb{I}(y_{i}^{\prime}=c)-f_{c}(\boldsymbol{x}_{i})\Bigr|$ (20)
$=\frac{1}{nk}\sum_{i=1}^{n}\sum_{c=1}^{k}\Bigl|\mathbb{I}(y_{i}^{\prime}=c)-f_{c}(\boldsymbol{x}_{i})\Bigr|$ (21)
which implies $I_{c,b}$ can be further divided into $n_{c,b}$ sub-partitions so that each sub-partition contains exactly one sample. Let $B^{\star}$ denotes this implied binning scheme. Then we find the lower bound of $\widehat{\mathrm{cwECE}}_{B^{\star}}(f)$. For the wrongly classified sample in $I_{c,b}$ where the prediction $\hat{y}_{i}\neq y_{i}^{\prime}$, we have
$\sum_{c=1}^{k}\bigl|\,\mathbb{I}(y_{i}^{\prime}=c)-f_{c}(\boldsymbol{x}_{i})\,\bigr|=2(1-f_{y_{i}^{\prime}}(\boldsymbol{x}_{i}))>1$ (22)
since only the true class contributes 11, $\sum_{c}f_{c}(\boldsymbol{x}_{i})=1$ and $f_{y_{i}^{\prime}}(\boldsymbol{x}_{i})<\frac{1}{2}$. For the correct classified sample in $I_{c,b}$, we have
$\sum_{c=1}^{k}\bigl|\,\mathbb{I}(y_{i}^{\prime}=c)-f_{c}(\boldsymbol{x}_{i})\,\bigr|<2(1-\frac{1}{k})$ (23)
since $1\geq f_{y_{i}^{\prime}}(\boldsymbol{x}_{i})>\frac{1}{k}$. Combine this equation with Eq. (21), then:
$\sup_{B\in\mathcal{B}}\,\widehat{\operatorname{cwECE}}_{B}(f)=\widehat{\mathrm{cwECE}}_{B^{\star}}(f)>\frac{1}{nk}\sum_{i=1}^{n}\mathbb{I}(\hat{y}_{i}\neq y_{i}^{\prime})=\frac{1}{k}\widehat{\operatorname{err}}(f).$
Then by applying Jensen’s inequality, we obtain the first inequality in this proposition. ∎
B.2 Proof of Prop. 4.6
Proof.
We begin with the expression for $\beta(\boldsymbol{x})$ from Eq. (11):
$\beta(\boldsymbol{x})=\mathbb{E}_{\tilde{\boldsymbol{x}}\sim P_{\mathcal{X}_{\tau}}}\left(\frac{\mathbb{I}\left[g(\boldsymbol{x})\geq g(\tilde{\boldsymbol{x}})\right]}{\mathbb{E}_{\boldsymbol{x}^{\prime}\sim\text{P}(\boldsymbol{x})}\mathbb{I}\left[g(\boldsymbol{x}^{\prime})\geq g(\tilde{\boldsymbol{x}})\right]}\right)=\mathbb{E}_{\tilde{\boldsymbol{x}}\sim P_{\mathcal{X}_{\tau}}}\left[\frac{\mathbb{I}[g(\boldsymbol{x})\geq g(\tilde{\boldsymbol{x}})]}{1-G(\tilde{\boldsymbol{x}})}\right$],
[…]
Next, we evaluate this expression by considering two different conditions for $\boldsymbol{x}$:
  • If $\boldsymbol{x}\in\mathcal{X}\setminus\mathcal{X}_{\tau}$:
[…]
    $\beta(\boldsymbol{x})=\int_{0}^{G(g^{-1}(\tau))}\frac{1}{1-t}dt,$
[…]
    $\int_{0}^{G(g^{-1}(\tau))}\frac{1}{1-t}dt=-\ln\left(1-G\left(g^{-1}\left(\tau\right)\right)\right).$
    Therefore, when $g(\boldsymbol{x})\geq\tau$, we obtain:
    $\beta(\boldsymbol{x})=-\ln(1-G(g^{-1}(\tau)).$
  • If $\boldsymbol{x}\in\mathcal{X}_{\tau}$:
[…]
    Similar to the first case, this results in:
    $\beta(\boldsymbol{x})=-\ln(1-G(\boldsymbol{x})).$
Combining the results from both conditions, we conclude that $\beta(\boldsymbol{x})$ is given by:
$\beta(\boldsymbol{x})=\begin{cases}-\ln(1-G(g^{-1}(\tau)))&\text{if }\boldsymbol{x}\in\mathcal{X}\setminus\mathcal{X}_{\tau},\\ -\ln(1-G(\boldsymbol{x}))&\text{if }\boldsymbol{x}\in\mathcal{X}_{\tau}.\end{cases}$
This completes the proof. ∎
Appendix C Gradient analysis
C.1 Focal loss
We consider the gradient of focal loss with respect to the true class predicted probability $p_{y^{\prime}}$. By applying the product rule, we obtain:
$\nabla_{p_{y^{\prime}}}\left[-(1-p_{y^{\prime}})^{\gamma}\log(p_{y^{\prime}})\right]$ $=\nabla_{p_{y^{\prime}}}\left[-(1-p_{y^{\prime}})^{\gamma}\right]\log(p_{y^{\prime}})+(1-p_{y^{\prime}})^{\gamma}\nabla_{p_{y^{\prime}}}\log(p_{y^{\prime}})$
$=\gamma(1-p_{y^{\prime}})^{\gamma-1}\log(p_{y^{\prime}})-(1-p_{y^{\prime}})^{\gamma}\frac{1}{p_{y^{\prime}}}.$
We can see that its gradient does shrink to 0 when $p_{y^{\prime}}\rightarrow 1$ (see Fig. 1).
(a) Focal loss
(b) Inverse focal
Figure 1: The gradient of loss $\ell$ with respect to $p_{y^{\prime}}$.
C.2 Inverse Focal loss
For inverse focal loss, similarly we can compute its gradient with respect to $p_{y^{\prime}}$:
$\nabla_{p_{y^{\prime}}}\left[-(1+p_{y^{\prime}})^{\gamma}\log(p_{y^{\prime}})\right]$ $=\nabla_{p_{y^{\prime}}}\left[-(1+p_{y^{\prime}})^{\gamma}\right]\log(p_{y^{\prime}})+(1+p_{y^{\prime}})^{\gamma}\nabla_{p_{y^{\prime}}}\log(p_{y^{\prime}})\$
$=\gamma(1+p_{y^{\prime}})^{\gamma-1}\log(p_{y^{\prime}})-(1+p_{y^{\prime}})^{\gamma}\frac{1}{p_{y^{\prime}}}$
where we can see that its gradient does not shrink to 0 when $p_{y^{\prime}}\rightarrow 1$ (see Fig. 1). However, we notice that when $\gamma$ is greater than 44, the gradient can be positive in some regions. In these cases, the inverse focal loss may not be suitable to optimize the model.
C.3 Dual Focal loss
(a) Weight
(b) Gradient
Figure 2: Illustration of DFL behavior on misclassified samples where $p_{y^{\prime}}<p_{j}$.
(c) Epoch 40
(d) Epoch 120
(e) Epoch 200
Figure 3: Evolution of the loss gradient $\partial\ell/\partial p_{y^{\prime}}$ with respect to model prediction probability $p_{y^{\prime}}$ for the same ResNet model evaluated on 20,000 CIFAR-10 training samples throughout training.
Dual Focal loss (DFL) Tao et al., (2023) generalizes Focal loss by incorporating both $p_{y^{\prime}}$ (the probability of the ground-truth class) and the probability of the most competitive alternative class, $p_{j}$. The intended formulation is
$\ell_{\text{DFL}}(\boldsymbol{p},\boldsymbol{y})\;=\;-\bigl(1-p_{y^{\prime}}+p_{j}\bigr)^{\gamma}\,\log p_{y^{\prime}}.$ (24)
Here, the weighting factor depends on the margin $p_{y^{\prime}}-p_{j}$, which quantifies the model’s uncertainty: samples with smaller margins (i.e., harder examples) are assigned larger weights and consequently incur higher penalties. In practice, however, the publicly released implementation diverges from (24). Specifically, the loss is computed as
$\ell_{\text{impl}}\big|_{p_{y^{\prime}}\geq p_{j}}$ $=-\bigl(1-p_{y^{\prime}}+p_{j}\bigr)^{\gamma}\,\log p_{y^{\prime}},$ (25)
$\ell_{\text{impl}}\big|_{p_{y^{\prime}}<p_{j}}$ $=-\bigl(1-p_{y^{\prime}}\bigr)^{\gamma}\,\log p_{y^{\prime}},$ (26)
which introduces a discontinuity at the tie point $p_{y^{\prime}}=p_{j}$. As $p_{y^{\prime}}$ falls below $p_{j}$, the competing probability is abruptly discarded and the formulation degenerates into the standard focal weighting, as shown in Fig. 2. This discontinuity results in the under-penalization of misclassified samples. More importantly, the realized variant acts in the opposite direction of the intended principle: by nullifying the strongest competitor precisely when $p_{y^{\prime}}<p_{j}$, it reduces the penalty for samples with large uncertainty. This contradiction resembles the effect observed in selective classification, where hard samples are down-weighted rather than emphasized.
C.4 $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ Loss
For the AU loss, we can compute the gradient with respect to the model output $\boldsymbol{p}_{i}=f_{\theta}(\boldsymbol{x}_{i})$ as follows:
$\nabla_{\boldsymbol{p}_{i}}\left(\widehat{\beta}_{i}\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})\right)$ $=(\nabla_{\boldsymbol{p}_{i}}\widehat{\beta}_{i})\,\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i})+\widehat{\beta}_{i}\,\nabla_{\boldsymbol{p}_{i}}\ell(\boldsymbol{p}_{i},\boldsymbol{y}_{i}),$ (27)
where the gradient of the sample weight $\widehat{\beta}_{i}$ is
$\nabla_{\boldsymbol{p}_{i}}\widehat{\beta}_{i}$ $=\begin{cases}\dfrac{1}{1-\widehat{G}_{\nu}(\boldsymbol{x}_{i})}\,\nabla_{\boldsymbol{p}_{i}}\widehat{G}_{\nu}(\boldsymbol{x}_{i}),&g(\boldsymbol{x}_{i})<\tau,\\[4.30554pt] 0,&g(\boldsymbol{x}_{i})\geq\tau,\end{cases}$
and the gradient of the smoothed bin-based CDF is
$\nabla_{\boldsymbol{p}_{i}}\widehat{G}_{\nu}(\boldsymbol{x}_{i})$ $=\frac{1}{n}\sum_{k=0}^{K-1}h_{k}\,\sigma^{\prime}\Big(\frac{g(\boldsymbol{x}_{i})-c_{k}}{\nu}\Big)\frac{1}{\nu}\,\nabla_{\boldsymbol{p}_{i}}g(\boldsymbol{x}_{i}),$
$\sigma^{\prime}(t)$ $=\sigma(t)(1-\sigma(t)).$
We can see that the gradient decomposes into two components: one corresponding to the weight $\widehat{\beta}_{i}$ (which depends on the smoothed CDF) and the other corresponding to the original loss $\ell$. For high-confidence samples ($g(\boldsymbol{x}_{i})\geq\tau$), the gradient reduces to the usual weighted loss term since $\nabla_{\boldsymbol{p}_{i}}\widehat{\beta}_{i}=0$.
Appendix D ADDITIONAL EXPERIMENTS
D.1 Experimental Setup
The experiments are conducted using NVIDIA A100 GPUs. All models are trained using stochastic gradient descent (SGD) with a momentum of 0.9. For CIFAR-10 and CIFAR-100, we use a mini-batch size of 128, an initial learning rate of 0.1, and an $\ell_{2}$ weight decay of 0.0005. The learning rate is decayed by a factor of 0.2 at epochs 60, 120, and 160, and the training is conducted for 200 epochs for CIFAR-10 and 250 epochs for CIFAR-100. For all transformer-based models trained on Tiny-ImageNet, we use a mini-batch size of 256, an initial learning rate of $1\times 10^{-4}$, a weight decay of 0.0005, and cosine learning rate scheduling for 90 epochs. For each loss function, we report the testing performance of the best model checkpoint selected based on the lowest validation loss after a warmup period (up to 40 epochs for CIFAR-10/100, 20 epochs for Tiny-ImageNet). Each configuration is run with three random seeds; we report the mean across runs.
Appendix E Model Selection
E.1 Baselines
Following Mukhoti et al., (2020); Tao et al., (2023), we perform cross-validation to select the optimal $\gamma$ value. We hold out 5,000 images from the training set as a validation set, train on the rest, and choose the best $\gamma$ based on validation performance i.e. loss (see Table 2 for the selected $\gamma$ values). Fig. 4 presents the performance on the validation set for Tiny-ImageNet data. For Inverse Focal Loss, Fig. 4a shows that $\gamma\geq 4$ consistently underperform, which should not be considered in practice. For KDE-XE, we use the regularization parameters recommended in the original work: $\lambda=0.3$ for CIFAR-10 and $\lambda=0.5$ for CIFAR-100.
(a) Inverse focal
(b) Dual focal
Figure 4: $\gamma$ selection for inverse focal loss using cross-validation on Tiny-Imagenet.
Table 2: Selected $\gamma$ values for Dual Focal and Inverse Focal loss across models and datasets.
Loss CIFAR-10 CIFAR-100 Tiny-ImageNet
ResNet50 ResNet110 WRN PreResNet56 ResNet50 ResNet110 WRN PreResNet56 ViT-Small Swin-Tiny
Dual 5.0 4.5 2.6 4.0 5.0 6.1 3.9 4.5 1.0 1.0
Inv. FL 2.0 3.0 1.5 2.0 1.0 1.5 1.0 1.0 1.0 1.0
E.2 $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ loss
This loss has two tunable parameters: the smooth parameter $\nu$ introduced by the CDF approximation and the threshold $\tau$ in Eq. (17). Smooth parameter. The smoothness parameter $\nu>0$ in the bin-based CDF approximation controls the steepness of the sigmoid function used to approximate the step function for each bin. As shown in Fig. 5, smaller values of $\nu$ produce a sharper transition, closely mimicking the discrete step function and providing a more accurate approximation of the true CDF, but may lead to large gradients and unstable training. Larger values of $\nu$ result in a flatter, smoother transition, which stabilizes gradient-based optimization but may over-smooth the CDF and reduce ranking precision in low-confidence regions. In practice, $\nu$ is typically chosen as a small fraction of the bin width or tuned via cross-validation, balancing the trade-off between approximation fidelity and differentiability for efficient optimization. We set this parameter to 0.1 for all models on CIFAR-10 and 0.01 for models except for WRN (0.1) on CIFAR-100. For transformer models, it is set to 0.1 for Swin-Tiny and 0.05 for ViT-Small.
(a) Histgram
(b) $\nu=0.01$
(c) $\nu=0.1$
(d) $\nu=1.0$
Figure 5: Effect of the smoothness parameter $\nu$ on the CDF approximation.
(a) Accuracy
(b) Gradient
Figure 6: Validation Performance of PreResNet56 on CIFAR-10 across low confidence coverage and $\nu$ Values
Coverage-Based Threshold $\tau$. Instead of using a fixed threshold, we determine $\tau$ dynamically based on a desired coverage level, which specifies the fraction of samples considered low-confidence. Under this scheme, the threshold is set such that the bottom $\kappa$ fraction of samples ranked by the CSF $g(\boldsymbol{x})$ are treated as low-confidence, while the remaining samples are classified as high-confidence and assigned uniform weights. This approach allows the loss to adapt to the distribution of confidence scores within each batch or dataset, ensuring that a consistent proportion of uncertain samples is emphasized for calibration, regardless of confidence values. Coverage-based thresholding provides a more flexible and robust mechanism for targeting low-confidence regions compared to a fixed $\tau$, particularly in settings where confidence distributions vary across datasets or model architectures. The selected low-confidence coverage for different models and datasets is summarized in Table 3. A subset of the selection results is presented in Table 6.
Table 3: Selected low confidence coverage for $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$ loss across models and datasets.
CIFAR-10 CIFAR-100 Tiny-ImageNet
ResNet50 ResNet110 WRN PreResNet56 ResNet50 ResNet110 WRN PreResNet56 ViT-Small Swin-Tiny
0.9 0.9 0.8 0.8 0.9 0.8 0.8 0.9 0.2 0.4
E.3 Additional Results
[…]
[Effect of post-hoc calibration method.] Although post-hoc calibration techniques such as MetaMis and MetaAcc are designed to enhance calibration by leveraging confidence information similar to our proposed method, they exhibit a noticeable trade-off between calibration quality and predictive performance, as indicated in Table 4. While both methods occasionally reduce ECE, the improvements are often offset by substantial accuracy degradation. For instance, accuracy drops exceeding 10% are observed in models like WRN on CIFAR-100. On Tiny-ImageNet, MetaAcc yields inconsistent results and, in some cases, negatively impacts calibration. These findings suggest that, despite their intended purpose, MetaMis and MetaAcc lack the robustness and reliability demonstrated by our proposed method.
Model XE FL-53 Inv. FL Dual KDE-XE AURC $\mathbf{\text{AU}_{P_{\mathcal{X}_{\tau}}}}$
CIFAR10
[…]
+TS 1.98 [94.5] (1.3) 3.61 [93.6] (1.0) 3.29 [94.4] (1.4) 1.98 [94.1] (1.1) 2.04 [94.5] (1.3) 1.10 [93.3] (1.2) 1.03 [92.3] (1.1)
+MetaMis 0.73 [84.7] 3.58 [86.6] 0.81 [84.3] 1.90 [85.5] 0.55 [84.3] 0.64 [87.1] 0.53 [87.6]
+MetaAcc 0.42 [77.5] 3.13 [74.9] 0.58 [78.7] 1.91 [72.0] 0.46 [76.9] 0.93 [90.0] 0.72 [89.9]
[…]
+TS 2.01 [94.5] (1.3) 3.93 [93.8] (1.0) 3.92 [93.9] (1.4) 2.45 [93.7] (1.0) 1.78 [94.7] (1.3) 1.24 [93.0] (1.2) 1.02 [92.4] (1.1)
+MetaMis 0.55 [84.5] 3.98 [87.6] 1.58 [86.7] 2.50 [86.1] 0.53 [86.0] 0.78 [87.6] 0.62 [89.0]
+MetaAcc 0.42 [78.1] 3.25 [74.2] 2.29 [88.1] 2.42 [77.2] 0.40 [78.0] 0.99 [91.6] 0.79 [90.5]
[…]
+TS 0.72 [95.9] (1.3) 5.13 [95.5] (1.0) 1.33 [96.0] (1.4) 5.27 [95.7] (1.1) 0.82 [95.8] (1.3) 0.74 [94.9] (1.2) 0.39 [95.1] (1.3)
+MetaMis 0.49 [82.6] 4.39 [86.2] 0.38 [82.6] 3.83 [84.8] 0.56 [82.5] 0.59 [87.3] 0.31 [86.7]
+MetaAcc 0.73 [71.6] 3.74 [69.2] 0.50 [72.4] 3.07 [66.4] 0.89 [71.0] 0.61 [82.1] 0.56 [75.4]
[…]
+TS 1.69 [94.0] (1.2) 3.06 [93.0] (0.9) 3.09 [94.1] (1.3) 2.61 [93.7] (1.0) 1.63 [94.1] (1.2) 1.18 [92.6] (1.2) 0.37 [92.8] (1.1)
+MetaMis 0.69 [85.9] 3.41 [88.7] 0.98 [84.8] 2.28 [86.0] 0.64 [85.3] 0.70 [87.8] 0.53 [89.2]
+MetaAcc 0.34 [75.9] 3.42 [75.8] 1.07 [82.7] 2.17 [68.2] 0.65 [81.9] 1.18 [91.5] 0.44 [89.6]
CIFAR100
[…]
+TS 3.10 [74.4] (1.2) 9.21 [76.4] (1.1) 4.15 [75.6] (1.4) 7.65 [75.5] (1.1) 3.12 [76.2] (1.3) 4.01 [75.4] (1.2) 1.78 [74.7] (1.1)
+MetaMis 1.99 [63.7] 5.36 [59.1] 1.51 [47.6] 4.56 [60.7] 1.20 [74.3] 2.78 [68.2] 1.19 [69.2]
+MetaAcc 2.13 [61.0] 3.30 [38.6] 0.95 [37.6] 2.62 [40.9] 0.68 [38.6] 3.38 [73.3] 1.40 [71.8]
[…]
+TS 3.46 [75.3] (1.2) 9.53 [77.5] (1.1) 4.56 [77.4] (1.4) 7.06 [76.3] (1.1) 3.03 [78.3] (1.4) 2.62 [74.3] (1.2) 1.66 [73.2] (1.1)
+MetaMis 2.04 [63.3] 5.16 [58.4] 1.40 [48.3] 4.21 [62.9] 0.98 [50.7] 2.90 [69.2] 1.43 [68.5]
+MetaAcc 2.12 [61.1] 3.24 [38.8] 0.93 [39.1] 2.70 [43.0] 0.61 [40.6] 3.52 [73.1] 1.57 [71.1]
[…]
+TS 5.00 [78.7] (1.3) 12.80 [79.1] (1.2) 4.99 [79.5] (1.4) 9.99 [78.4] (1.2) 4.71 [79.6] (1.4) 2.98 [76.5] (1.2) 6.59 [79.8] (1.3)
+MetaMis 1.01 [54.4] 5.45 [56.8] 0.70 [46.4] 3.93 [57.8] 0.66 [46.9] 1.29 [69.5] 0.44 [47.8]
+MetaAcc 1.11 [48.2] 2.60 [34.3] 0.50 [34.9] 1.75 [38.0] 0.45 [37.0] 1.98 [74.0] 0.24 [34.6]
[…]
+TS 5.13 [74.3] (1.1) 1.28 [72.9] (0.9) 7.08 [74.2] (1.1) 1.78 [73.0] (0.9) 5.27 [74.7] (1.1) 3.10 [71.9] (1.1) 2.13 [73.9] (1.0)
+MetaMis 3.56 [69.3] 1.62 [67.8] 5.22 [69.2] 1.04 [70.4] 3.79 [70.0] 2.17 [68.2] 1.23 [70.9]
+MetaAcc 4.00 [70.6] 1.43 [68.6] 6.03 [71.8] 1.19 [70.9] 4.15 [71.0] 2.72 [71.2] 1.70 [73.0]
Tiny-ImageNet
[…]
+TS 3.69 [82.9] (1.3) 8.43 [81.4] (1.1) 3.92 [82.9] (1.3) 4.25 [82.1] (1.2) 5.03 [83.3] (1.2) 6.67 [83.1] (1.2) 7.37 [83.4] (1.2)
+MetaMis 3.23 [76.9] 8.71 [76.7] 2.86 [74.6] 3.99 [75.5] 4.63 [77.8] 5.68 [77.6] 7.04 [78.3]
+MetaAcc 3.39 [80.5] 8.38 [64.3] 2.80 [70.7] 3.89 [70.3] 4.94 [81.3] 6.47 [82.1] 7.28 [82.6]
[…]
+TS 5.55 [85.4] (1.3) 13.50 [84.3] (1.1) 3.56 [85.6] (1.3) 8.91 [85.4] (1.2) 5.55 [85.4] (1.3) 8.65 [82.4] (1.2) 7.53 [84.4] (1.2)
+MetaMis 5.12 [81.7] 13.28 [80.7] 3.12 [81.2] 8.53 [81.8] 5.07 [81.3] 7.67 [77.3] 6.81 [79.5]
+MetaAcc 5.37 [84.2] 13.30 [82.5] 3.29 [84.3] 8.70 [84.2] 5.36 [84.2] 8.40 [81.3] 7.35 [83.3]
Table 4: Mean ECEa $\times 10^{2}$ ($\downarrow$) for original models, original models, and after applying post-hoc calibration methods including TS, MetaMis, and MetaAcc. Accuracy (%) shown in brackets. For TS, temperature values are included in parentheses. For TS, the temperature $T$ is selected by minimizing ECE on the training set.